OpenStudy (anonymous):

how do i write this as an expression? cos(pi/5)cos(pi/7)+sin(pi/5)sin(pi/7)

OpenStudy (freckles):

recall \[\cos(x)\cos(y)+\sin(x)\sin(y)=\cos(x-y)\]

OpenStudy (freckles):

though it is already an expression I think you meant to simplify the given expression

Nnesha (nnesha):

it would be easy if you change radians to degree hahah i like degrees o^_^o

OpenStudy (anonymous):

how do i write the expression as either the sine, cosine, or tangent of a single angle

Nnesha (nnesha):

recall \[\rm cos(x)\cos(y)+\sin(x)\sin(y)=\color{reD}{\cos(x-y)}\] yes you have to find red part

Nnesha (nnesha):

cos pi /5 = ? cos pi /7 = ? pi/ 5 = x pi /7 ] y

OpenStudy (anonymous):

thank you

Nnesha (nnesha):

np :-) let us know what you got :-)